What is the pH of a solution containing 0.30 M HCOOH and 0.52 M HCOOK?

How do i get the pKa?? I know its 3.77 but how??

Help!

Edited; Nvm. figured it out, thanks
Answer #1

According to Henderson's Equation , pH = pK_{a}+ log\frac{\left [ salt \right ]}{\left [ acid \right ]}

                                                         pH = pK_{a}+ log\frac{\left [HCOOK ]}{\left [ HCOOH ]}

Given [HCOOK] = 0.52 M

          [ HCOOH] = 0.30 M

The value of acid dissociation constant for formic acid from the standard list is, Ka =1.7x10 -4

So pKa = - log Ka

             = - log 1.7x10 -4

             = 3.77

Plug these values in Henderson's equation we have pH = 3.77+ log\frac{\left 0.52}{\left 0.30}

                                                                                 pH = 3.77+ 0.24

                                                                                 pH =4.01

Therefore the pH of the resulting solution is 4.01